Questions AMOXICILLIN TEVA Y TUTOP lauk
Answer #1

Ans) a) Recommended reconstitution: 80-90 ml as the medication is available in powder form.

There is always marking on the bottle so that the quantity & dose is appropriate.

b) Recommend dilueDi is boiled cooled water.

c) Dosage strength of recommended solution:

125 mg = 5 ml

0.4g = 0.4 × 1000 = 400 mg is the dosage Strength.

d. 125 mg = 5 ml

400 mg = x

x = 400 × 5/ 125 = 16 ml

= 2000/125 = 16